OpenStudy (anonymous):

can anyone help me out with this please :-) :-P :-D "_" 0^_^0 ...........?

OpenStudy (anonymous):

The question is A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (austinl):

Because it is an absolute value we can break it into two parts, \(x-0.525=0.01\) & \(x+0.525=0.01\) Solve those for x and boom, there you have it.